# Hamiltonian stationary Lagrangian surfaces with harmonic mean curvature in complex space forms

### v4

@@ Abstract.
In this paper, we deal with Hamiltonian stationary Lagrangian surfaces in complex space forms whose mean curvature is a harmonic function. We prove that if the mean curvature is non-zero constant, then the second fundamental form is parallel. In the non-constant mean curvature case, under the condition of constant Gaussian curvature, we give an explicit description of all such surfaces.
@@ 1. Introduction
(1.1) ${\rm div}(JH)=0$
on $M$ (cf. [9]), where div is the divergence on $M$.
In this paper, we deal with Hamiltonian stationary Lagrangian surfaces in complex space forms. First, we prove that if the mean curvature is non-zero constant, then the second fundamental form is parallel (Proposition 3.1). Next, we investigate the case where the mean curvature is a non-constant harmonic function. Our purpose is to give an explicit description of all such surfaces under the condition of constant Gaussian curvature (Theorem 4.1).
@@ 2. Preliminaries
Let $M$ be a Lagrangian submanifold of $\tilde{M}^{n}(4\epsilon)$. We denote the Levi-Civita connections on $M^{n}$ and $\tilde{M}^{n}(4\epsilon)$ by $\nabla$ and $\tilde{\nabla}$, respectively. The Gauss and Weingarten formulas are respectively given by
$\tilde{\nabla}_{X}Y=\nabla_{X}Y+h(X,Y),\quad\tilde{\nabla}_{X}\xi=-A_{\xi}X+D_{X}\xi$
for tangent vector fields $X$, $Y$ and normal vector field $\xi$, where $h,A$ and $D$ are the second fundamental form, the shape operator and the normal connection. The mean curvature vector field $H$ is defined by $H=(1/n){\rm trace}\,h.$ The function $|H|$ is called the mean curvature. We have (cf. [7])
@@
Denote by $R$ the Riemann curvature tensor of $\nabla$. Then the equations of Gauss and Codazzi are respectively given by
@@
3. Hamiltonian stationary Lagrangian surfaces
Let $M$ be a Hamiltonian stationary Lagrangian surface in $\tilde{M}^{2}(4\epsilon)$, where $\epsilon\in\{-1,0,1\}$. Suppose that $H\neq 0$ everywhere. Denote by $K$ the Gaussian curvature of $M$. Let $\{e_{1},e_{2}\}$ be a local orthonormal basis of $M$ such that $Je_{1}$ is parallel to $H$. It follows from (2.2) that second fundamental form takes the form
@@ 3. Hamiltonian stationary Lagrangian surfaces
for some functions $a$, $b$ and $c$. Putting $\omega_{i}^{j}(e_{k})=\left<\nabla_{e_{k}}e_{i},e_{j}\right$>, by (2.1) and (LABEL:sf) we have
@@
Therefore, the equation (2.4) of Codazzi implies
@@
The Hamiltonian stationary condition (1.1) is equivalent to
@@
Therefore, there exists a local coordinate system $\{u,v\}$ such that $e_{1}=a\partial_{u}$, $e_{2}=a\partial_{v}$. Hence, the metric tensor is given by
@@
Equations (3.2) and (3.3) are rewritten as
@@
Put $\delta=\epsilon-K$. Then the equation (2.3) of Gauss and (LABEL:sf) yield that
@@
Let $M$ be a Hamiltonian stationary Lagrangian surface in $\tilde{M}^{2}(4\epsilon)$. If $|H|$ is non-zero constant, then $\bar{\nabla}h=0$.
Proof. From (3.5) and (3.6), it follows that if $|H|=|a|/2$ is non-zero constant, then $\omega_{1}^{2}=0$. Hence we have $K=0$. Differentiating (3.12) along $e_{i}$ $(i=1,2)$ yields $4be_{i}b+(4c-a)e_{i}c=0$, which together with (3.2) and (3.3), shows that $b$ and $c$ are constant. Thus, using $\omega_{1}^{2}=0$, (2.1) and (LABEL:sf), we obtain $\bar{\nabla}h=0$.
@@
An explicit description of all Lagrangian surfaces with $\bar{\nabla}h=0$ in $\tilde{M}^{2}(4\epsilon)$ has been obtained in Theorem 7.2 in [1], Theorem A.1 and Theorem A.2 in [4].
4. Main theorem
Let ${\mathbb{C}}^{3}_{1}$ be the complex 33-space endowed with the inner product
$\left<(z_{1},z_{2},z_{3}),(w_{1},w_{2},w_{3})\right>={\rm Re}(-z_{1}\bar{w}_{1}+z_{2}\bar{w}_{2}+z_{3}\bar{w}_{3}).$
Put $H_{1}^{5}(-1)=\{z\in{\mathbb{C}}^{3}_{1}:\left<z,z\right>=-1\}$. The map $\Pi:H_{1}^{5}(-1)\rightarrow{\mathbb{C}}H^{2}(-4)$ defined by $z\mapsto z\cdot{\mathbb{C}}^{*}$ is referred to as the Hopf fibration.
The main result of this paper is the following classification theorem.
@@ 4. Main theorem
Let $M$ be a Hamiltonian stationary Lagrangian surface in a 22-dimensional complex space form $\tilde{M}^{2}(4\epsilon)$, where $\epsilon\in\{-1,0,1\}$. Suppose that $H$ is nowhere vanishing. If $M$ has non-constant harmonic mean curvature and constant Gaussian curvature $K$, then $K=\epsilon=-1$ and $M$ is locally congruent to the image of $\Pi\circ\phi$, where $\Pi:H_{1}^{5}(-1)\rightarrow{\mathbb{C}}H^{2}(-4)$ is the Hopf fibration and $\phi:M\rightarrow H_{1}^{5}(-1)\subset{\mathbb{C}}^{3}_{1}$ is given by one of the following immersions::
@@
Proof. We shall use the same notation as in Section 3. Suppose that $|H|$ is a harmonic function on $M$. Then by (3.7) we have
@@
for some function $\theta$ on $M$. Substituting (4.2) into (4.1), we have
@@
It follows (4.3) from (4.4) that $\theta_{u}=\theta_{v}=0$, that is, $\theta$ is constant. Solving (4.2), we conclude that up to translations, $a$ is given by
@@
If $c_{v}\neq 0$ on an open subset in $\mathcal{U}$, then $a=c$ on that open subset. From (3.12) we see that $a$ is constant, which contradicts our assumption. Hence we have $c_{v}=0$ on $\mathcal{U}$. Thus, by (4.7) we get
@@
Case (i.1): $c=0$ on an open subset $\mathcal{U}_{1}\subset\mathcal{U}$. In this case, from (3.12) we have $K=\epsilon=-1$. It follows from (4.6) that $a_{u}=0$ on $\mathcal{U}_{1}$. Hence, using (4.5) and K=1K=-1 yields that $a^{2}=v^{2}$. Applying the coordinate transformation: $y=-\int v^{-1}dv$, we see that the metric tensor (3.7) becomes
@@
Rewriting $u$ with $x$ and according to [3, page 3475], we conclude that $\mathcal{U}_{1}$ is congruent to the Lagrangian surface obtained from (1).
@@
If $a\neq 2c$ on an open set in $\mathcal{U}_{2}$, then $a=6c$ and it follows from (3.12) that $a$ is constant, which contradicts our assumption. Hence $a=2c$, which together with (3.12) implies $K=\epsilon=-1$. Thus, by (4.5) we obtain $a^{2}=u^{2}$. Applying the coordinate transformation: $x=(u+v)/2$ and y=(uv)/2y=(u-v)/2, we see that the metric tensor (3.7) becomes
@@
Case (ii): $b\neq 0$ on an open subset $\mathcal{V}$. We put $A=a_{u}$ and $B=a_{v}$, which are constants. Differentiating (3.12) with respect to $u$ and $v$, we obtain
@@
(4.12) $4abc_{u}-a(a-4c)c_{v}=4b(3c-a)A-(12b^{2}-ac)B,$
(4.13) $a(a-4c)c_{u}+4abc_{v}=(12b^{2}-ac)A+12bcB.$
Solving (4.12) and (4.13) for $c_{u}$ and $c_{v}$ and using (3.12), we have
(4.14) $c_{u}=\frac{Af_{1}+Bf_{2}}{a(a^{2}+8\delta)},$
(4.15) $c_{v}=\frac{Af_{3}+Bf_{4}}{a(a^{2}+8\delta)},$
@@
(4.16) $b_{u}=$ $\frac{Ag_{1}+Bg_{2}}{a(a^{2}+8\delta)},$
(4.17) $b_{v}=$ $\frac{Ag_{3}+Bg_{4}}{a(a^{2}+8\delta)},$
@@
Differentiating (4.14) and (4.15) with respect to $v$ and $u$ respectively, we have
(4.18) $c_{uv}=\frac{(Af_{1}+Bf_{2})_{v}a(a^{2}+8\delta)-(Af_{1}+Bf_{2})(3a^{2}+8\delta)B}{a^{2}(a^{2}+8\delta)^{2}},$
(4.19) $c_{vu}=\frac{(Af_{3}+Bf_{4})_{u}a(a^{2}+8\delta)-(Af_{3}+Bf_{4})(3a^{2}+8\delta)A}{a^{2}(a^{2}+8\delta)^{2}}.$
By a straightforward but tedious computation, using (4.14)-(4.17) and substituting (3.12) into the numerators of (4.18) and (4.19), we get
(4.20) $c_{uv}=\frac{A^{2}h_{1}+ABh_{2}+B^{2}h_{3}}{a^{2}(a^{2}+8\delta)^{2}},$
(4.21) $c_{vu}=\frac{A^{2}h_{4}+ABh_{5}+B^{2}h_{6}}{a^{2}(a^{2}+8\delta)^{2}},$
@@
Since $c_{uv}-c_{vu}$=0 holds, from (4.20), (4.21), it follows that
(4.22) $4b(A^{2}k_{1}+ABk_{2}+B^{2}k_{3})=0,$
@@
By a straightforward but tedious computation, using (4.14)-(4.17), we get
(4.23) $(A^{2}k_{1}+ABk_{2}+B^{2}k_{3})_{u}$ $=\frac{4(A^{3}P_{1}+A^{2}BP_{2}+AB^{2}P_{3}+B^{3}P_{4})}{a^{2}+8\delta},$
(4.24) $(A^{2}k_{1}+ABk_{2}+B^{2}k_{3})_{v}$ $=\frac{8(A^{3}P_{5}+A^{2}BP_{6}+AB^{2}P_{7}+B^{3}P_{8})}{a^{2}+8\delta},$
@@
Using a computer algebra system, we check that the resultant of the left-hand sides of (4.25) and (4.26) with respect to AA is, up to a constant factor,
@@
Combining (3.12) and (4.27), we find that (4.27) can be simplified to
(4.28) $B^{9}\delta^{3}(a^{2}+8\delta)^{7}(a^{2}+24\delta)^{2}(5a^{6}+60a^{4}\delta+180a^{2}\delta^{2}+288\delta^{3}),$
which must vanish on $\mathcal{V}$. Taking into account $\nabla a\neq 0$, we have
@@
From (3.12) we see that $c\neq 0$. Changing the sign of $e_{1}$ if necessary, we may assume that $c>0$. Put
@@
We make the following change of the basis:
@@
Therefore, according to [5, Theorem 4.1], there exists a local coordinate system $\{x,y\}$ such that $\partial_{x}=f\tilde{e}_{1}$ and $\partial_{y}=k\tilde{e}_{2}$, where $f$ and $k$ satisfy
@@
where $\lambda$ and $m$ are positive real numbers.
We note that $f=mk$. By the first equation of (4.33) we have
@@
Since $a_{u}$ and $a_{v}$ are constant, the solutions (4.37) and (4.38) are excluded. Considering also that $a_{u}^{2}+a_{v}^{2}=1$ and $f\neq k$, we see that $f$ and $g$ are given by (4.39) with $m\neq 1$. In [6, Section 7], it is proved that the corresponding surface is congruent to the Lagrangian surface obtained from case (2) with $m\neq 1$. This completes the proof.

### v5

@@ Abstract.
In this paper, we study Hamiltonian stationary Lagrangian surfaces in complex space forms. We first show that when the mean curvature is a non-zero constant, the second fundamental form is parallel. We then consider the case in which the mean curvature is a non-constant harmonic function. Under the additional assumption that the Gaussian curvature is constant, we obtain a complete classification of such Lagrangian surfaces.
@@ 1. Introduction
(1.1) ${\rm div}(JH)=0,$
where $H$ denotes the mean curvature vector field of $M$.
In this paper, we investigate Hamiltonian stationary Lagrangian surfaces in complex space forms. When the mean curvature is a non-zero constant, we show that the second fundamental form is parallel. Such Lagrangian surfaces have been completely classified in Theorem 7.2 of [1], Theorems A.1 and A.2 of [4]. We then turn to the case in which the mean curvature is a non-constant harmonic function. Assuming that the Gaussian curvature is constant, we prove that the ambient space must be the complex hyperbolic plane and the Gaussian curvature necessarily takes a specific negative value. This phenomenon leads to a complete classification.
@@ 2. Preliminaries
Let $M$ be a Lagrangian submanifold of $\tilde{M}^{n}(4\epsilon)$. We denote the Levi-Civita connections on $M^{n}$ and $\tilde{M}^{n}(4\epsilon)$ by $\nabla$ and $\tilde{\nabla}$, respectively. The Gauss and Weingarten formulas are given respectively by
$\tilde{\nabla}_{X}Y=\nabla_{X}Y+h(X,Y),\quad\tilde{\nabla}_{X}\xi=-S_{\xi}X+D_{X}\xi$
for tangent vector fields $X$, $Y$ and normal vector field $\xi$, where $h$, $S$ and $D$ are the second fundamental form, the shape operator and the normal connection. The mean curvature vector field $H$ is defined by $H=(1/n){\rm trace}\,h.$ The function $|H|$ is called the mean curvature. We have (cf. [7])
@@
Denote by $R$ the Riemann curvature tensor of $\nabla$. Then the equations of Gauss and Codazzi are given respectively by
@@
3. Hamiltonian stationary Lagrangian surfaces with harmonic mean curvature
Let $M$ be a Lagrangian surface in $\tilde{M}^{2}(4\epsilon)$, where $\epsilon\in\{-1,0,1\}$. Suppose that $H\neq 0$ everywhere. Denote by $K$ the Gaussian curvature of $M$. Let $\{e_{1},e_{2}\}$ be a local orthonormal frame on $M$ such that $Je_{1}$ is parallel to $H$. It follows from (2.2) that the second fundamental form takes the form
@@ 3. Hamiltonian stationary Lagrangian surfaces with harmonic mean curvature
for some functions $a$, $b$ and $c$.
Putting $\omega_{i}^{j}(e_{k})=\left<\nabla_{e_{k}}e_{i},e_{j}\right$>, and using (2.1) and (LABEL:sf), we have
@@
Therefore, the Codazzi equation (2.4) implies
@@
Assume that $M$ is Hamiltonian stationary. Then, by (1.1) we have
@@
Therefore, there exists a local coordinate system $\{u,v\}$ such that $e_{1}=a\partial_{u}$ and $e_{2}=a\partial_{v}$. Hence, the metric tensor is given by
@@
It follows from (3.8) that (3.2) and (3.3) can be rewritten as
@@
Put $\delta=\epsilon-K$. Then the Gauss equation (2.3) together with (LABEL:sf) yields
@@
In the case where the mean curvature is constant, we have
@@
Let $M$ be a Hamiltonian stationary Lagrangian surface in $\tilde{M}^{2}(4\epsilon)$. If $|H|$ is a non-zero constant, then $\bar{\nabla}h=0$.
Proof. From (3.8), it follows that if $|H|=|a|/2$ is a non-zero constant, then $\omega_{1}^{2}=0$, and hence $K=0$. Equations (3.2) and (3.3) reduce to
(3.13) $c_{u}-b_{v}=c_{v}+b_{u}=0.$
Since $\delta=\epsilon$, differentiating (3.12) with respect to $u$ and $v$, and using (3.13), we have
(3.14) $\begin{pmatrix}4b&4c-a\\ a-4c&4b\\ \end{pmatrix}\begin{pmatrix}b_{u}\\ c_{u}\end{pmatrix}=\begin{pmatrix}0\\ 0\end{pmatrix}$.
If $16b^{2}+(a-4c)^{2}\neq 0$ on an open subset, then by (3.14) we obtain $b_{u}=c_{u}=0$. Thus, it follows from (3.13) that $b$ and $c$ are constant. Otherwise, $b=0$ and $c$ is constant. In both cases, using (2.1) and (LABEL:sf), we obtain $\bar{\nabla}h=0$. ∎
@@
An explicit description of all Lagrangian surfaces with $\bar{\nabla}h=0$ in $\tilde{M}^{2}(4\epsilon)$ has been obtained in Theorem 7.2 of [1], Theorems A.1 and A.2 of [4].
Our next step is to investigate the case of non-constant mean curvature. The main result of this paper is the following.
@@
Let $M$ be a Hamiltonian stationary Lagrangian surface in a complex space form $\tilde{M}^{2}(4\epsilon)$, where $\epsilon\in\{-1,0,1\}$. Suppose that $H$ is nowhere vanishing. If $|H|$ is a non-constant harmonic function and $K$ is constant, then $K=\epsilon=-1$ and $M$ is locally congruent to the image of $\Pi\circ\phi$, where $\Pi:H_{1}^{5}(-1)\rightarrow{\mathbb{C}}H^{2}(-4)$ is the Hopf fibration and $\phi:M\rightarrow H_{1}^{5}(-1)\subset{\mathbb{C}}^{3}_{1}$ is given by one of the following immersions::
@@
Proof. Suppose that $|H|$ is a harmonic function on $M$. Then by (3.7) we have
@@
for some function $\theta$ on $M$. Substituting (3.16) into (3.15), we get
@@
It follows from (3.17) and (3.18) that $\theta_{u}=\theta_{v}=0$, that is, $\theta$ is constant. Solving (3.16), we conclude that up to translations, $a$ is given by
@@
If $c_{v}\neq 0$ on an open subset in $\mathcal{U}$, then $a=c$. From (3.12) we see that $a$ is constant, which contradicts our assumption. Hence we have $c_{v}=0$ on $\mathcal{U}$. Thus, by (3.21) we get
@@
Case (i.1): $c=0$ on an open subset $\mathcal{U}_{1}\subset\mathcal{U}$. In this case, from (3.12) we have $K=\epsilon=-1$. It follows from (3.20) that $a_{u}=0$ on $\mathcal{U}_{1}$. Hence, using (3.19) and K=1K=-1, we obtain $a^{2}=v^{2}$. Applying the coordinate transformation $y=-\int v^{-1}dv$, we see that the metric tensor (3.7) becomes
@@
We rewrite $u$ as $x$. According to [3, page 3475], we conclude that $\mathcal{U}_{1}$ is congruent to the Lagrangian surface obtained from (1).
@@
If $a\neq 2c$ on an open subset of $\mathcal{U}_{2}$, then $a=6c$, and it follows from (3.12) that $a$ is constant, which contradicts our assumption. Hence $a=2c$, which together with (3.12) implies $K=\epsilon=-1$. Thus, by (3.19) we obtain $a^{2}=u^{2}$. Applying the coordinate transformation $x=(u+v)/2$ and y=(uv)/2y=(u-v)/2, we see that the metric tensor (3.7) becomes
@@
Case (ii): $b\neq 0$ on an open subset $\mathcal{V}$. We put $A=a_{u}$ and $B=a_{v}$, which are constant. Differentiating (3.12) with respect to $u$ and $v$, we obtain
@@
(3.26) $a\{4bc_{u}-(a-4c)c_{v}\}=4b(3c-a)A-(12b^{2}-ac)B,$
(3.27) $a\{(a-4c)c_{u}+4bc_{v}\}=(12b^{2}-ac)A+12bcB.$
Solving (3.26) and (3.27) for $c_{u}$ and $c_{v}$, we have
(3.28) $c_{u}=\frac{Af_{1}+Bf_{2}}{a(a^{2}+16b^{2}-8ac+16c^{2})},$
(3.29) $c_{v}=\frac{Af_{3}+Bf_{4}}{a(a^{2}+16b^{2}-8ac+16c^{2})},$
@@
(3.34) $b_{u}=$ $\frac{Ag_{1}+Bg_{2}}{a(a^{2}+16b^{2}-8ac+16c^{2})},$
(3.35) $b_{v}=$ $\frac{Ag_{3}+Bg_{4}}{a(a^{2}+16b^{2}-8ac+16c^{2})},$
@@
Using (3.28)-(3.35), we differentiate (3.28) and (3.29) with respect to $v$ and $u$, respectively. A computation with a computer algebra system yields
(3.36) $c_{uv}=\frac{A^{2}h_{1}+ABh_{2}+B^{2}h_{3}}{a^{2}(a^{2}+16b^{2}-8ac+16c^{2})^{2}},$
(3.37) $c_{vu}=\frac{A^{2}h_{4}+ABh_{5}+B^{2}h_{6}}{a^{2}(a^{2}+16b^{2}-8ac+16c^{2})^{2}},$
@@
Subtracting (3.37) from (3.36), we have
(3.38) $c_{uv}-c_{vu}=\frac{4b(A^{2}k_{1}+ABk_{2}+B^{2}k_{3})}{a^{2}(a^{2}+16b^{2}-8ac+16c^{2})^{2}},$
@@
[…]
(3.39) $A^{2}k_{1}+ABk_{2}+B^{2}k_{3}=0.$
Differentiating the left hand side of (3.39) and using (3.28)-(3.35), with the aid of a computer algebra system, we get
(3.40) $(A^{2}k_{1}+ABk_{2}+B^{2}k_{3})_{u}$ $=\frac{4(A^{3}P_{1}+A^{2}BP_{2}+AB^{2}P_{3}+B^{3}P_{4})}{a^{2}+16b^{2}-8ac+16c^{2}},$
(3.41) $(A^{2}k_{1}+ABk_{2}+B^{2}k_{3})_{v}$ $=\frac{8(A^{3}P_{5}+A^{2}BP_{6}+AB^{2}P_{7}+B^{3}P_{8})}{a^{2}+16b^{2}-8ac+16c^{2}},$
@@
Using a computer algebra system, we find that the resultant of the left-hand sides of (3.42) and (3.43) with respect to AA is, up to a non-zero constant factor,
@@
Combining (3.12) and (3.44), we find that (3.44) simplifies to
$B^{9}\delta^{3}(a^{2}+8\delta)^{7}(a^{2}+24\delta)^{2}(5a^{6}+60a^{4}\delta+180a^{2}\delta^{2}+288\delta^{3}),$
which must vanish on $\mathcal{V}$. Taking into account that $\nabla a\neq 0$, we have
@@
From (3.12) and $b\neq 0$, we see that $c\neq 0$. Changing the sign of $e_{1}$ if necessary, we may assume that $c>0$. Put
@@
We make the following change of basis:
@@
According to [5, Theorem 4.1], there exists a local coordinate system $\{x,y\}$ such that $\partial_{x}=f\tilde{e}_{1}$ and $\partial_{y}=k\tilde{e}_{2}$, where $f$ and $k$ satisfy
@@
where $\lambda$ and $m$ are positive real numbers. We note that $f=mk$. By the first equation of (3.49) we have
@@
Since $a_{u}$ and $a_{v}$ are constant, the solutions (3.53) and (3.54) are excluded. Considering also that $a_{u}^{2}+a_{v}^{2}=1$ and $f\neq k$, we see that $f$ and $g$ are given by (3.55) with $m\neq 1$. In [6, Section 7], it is proved that the corresponding surface is congruent to the Lagrangian surface obtained from case (2) with $m\neq 1$. ∎